What is the payback period for the following project:

T=0

T=1

T=2

T=3

Cashflow

-$180,000

$93,600

$64,800

$81,600
Answer #1

Cumulative cash flow for year 0 = -180,000

Cumulative cash flow for year 1 = -180,000 + 93,600 = -86,400

Cumulative cash flow for year 2 = -86,400 + 64,800 = -21,600

Cumulative cash flow for year 3 = -21,600 + 81,600 = 60,000

21,600 / 81,600 = 0.26

Payback period = 2 + 0.26 = 2.26 years